### v1

@@ 6.1 Significance for Healthcare Management
Our open-source infrastructure for conducting network analysis on medical claims data opens the door for exciting advancements in the field of healthcare management. By leveraging network analysis techniques, we can uncover intricate patterns and relationships within healthcare delivery systems that were previously hidden. This is crucial for several reasons:
  • Optimization of Care Coordination: Understanding referral patterns and physician collaborations can help identify inefficiencies and bottlenecks in patient flow, leading to more effective care coordination.
  • Resource Allocation: Network analysis can inform strategic decisions regarding resource distribution, ensuring that healthcare facilities and services are optimally positioned to meet patient needs.
  • Policy Development: Insights from network structures can guide policymakers in developing initiatives that foster more robust and integrated healthcare systems.
@@ 6.3 Exploration Prototype
To further the impact of our research, we provide an open-access dataset and analysis suite, apparent [^5] Our prototype can be found at: https://apparent.topology.rocks/ that combines medical claims data with various demographic and socioeconomic factors. This provides access to network features and abundant metadata, along with prototypes for exploring/visualizing user queries. Moreover, we provide a Python repository [^6] Our codebase is maintained at: https://github.com/aidos-lab/apparent for interacting directly with the dataset that supports:
  • Reproducibility Package (Networks and Provided Features)
  • Visualizing Networks and Feature Distributions
  • Customizable Ricci Curvature Computations
By offering access to curvature features and other network metrics, we enable both technical and non-technical researchers to delve into the data, fostering a collaborative environment for discovering new insights.
@@ 6.4 Limitations
  • Computational Tractability: Calculating $\kappa_{\text{OR}}$ on massive networks is computationally intensive, limiting its scalability. Thus, we recommend users to prioritize $\kappa_{\text{FR}}$ for studying large scale Physician Referral Networks.
  • Network Specifications: We only build and analyze unweighted, undirected networks, which may overlook the complexities of real-world healthcare interactions. We hope to extend our analyses to more complex network constructions in the future.
  • Feature Completeness: Our prototype may lack some critical features necessary for a comprehensive understanding of healthcare management. This prototype serves as a foundation for future enhancements, and can easily be updated as necessary.
@@ 6.5 Future Work
  • Network Evolution: Studying how healthcare networks evolve over time to identify trends and predict future changes.
  • Thorough Data Analysis: Conducting a more detailed analysis of existing data to derive actionable insights that can influence policy and management decisions.
  • Regional Characteristics: Understanding the distinctive features of regional healthcare networks and how they relate to specific policies and practices.
  • Edge Prediction: Developing methods for predicting new referrals that could alleviate bottlenecks in patient flow, utilizing negative curvature as a predictive metric.
@@ Acknowledgements
We would like to acknowledge the authors and developers of the Datasette package, whose work has been instrumental in the development of our open-source prototype. Their tool has enabled us to efficiently manage, explore, and visualize our dataset, significantly enhancing our ability to analyze and share our findings. We also thank members of the CARE-CONNECT team. Research reported in this publication was supported by the National Heart, Lung, And Blood Institute of the National Institutes of Health under Award Number R01HL167816 as well as by the Helmholtz Association Initiative and Networking Fund on the HAICORE@KIT partition. The content is solely the responsibility of the authors and does not necessarily represent the official views of the National Institutes of Health.

### v2

@@ 6.1 Significance for Healthcare Management
Our open-source infrastructure enables network analysis on medical claims data. This opens the door for exciting advancements in healthcare management. By leveraging network analysis techniques, we can uncover intricate patterns and relationships within healthcare delivery systems that were previously hidden. This is crucial for several reasons:
  • Optimization of Care Coordination. Understanding referral patterns and physician collaborations can help identify inefficiencies and bottlenecks in patient flow, leading to more effective care coordination.
  • Resource Allocation. Network analysis can inform strategic decisions regarding resource distribution, ensuring that healthcare facilities and services are optimally positioned to meet patient needs.
  • Policy Development. Insights from network structures can guide policymakers in developing initiatives that foster more robust and integrated healthcare systems.
@@ 6.3 Exploration Prototype
To further the impact of our research, we provide an open-access dataset and analysis suite, apparent that combines medical claims data with various demographic and socioeconomic factors. This provides access to network features and abundant metadata, along with prototypes for exploring/visualizing user queries. Additionally, we provide a Python repository [^2] Our codebase is maintained at: https://github.com/aidos-lab/apparent that allows users to interact directly with the dataset. This repository supports building, customizing, and reproducing Physician Referral Networks and associated structural features, as well as visualizing networks and feature distributions. By offering access to curvature features and other network metrics, we empower both technical and non-technical researchers to explore the data, fostering a collaborative environment for discovering new insights.
@@ 6.4 Limitations
  • Computational Tractability. Calculating $\kappa_{\text{OR}}$ on massive networks is computationally intensive, limiting its scalability. Thus, we recommend users to prioritize $\kappa_{\text{FR}}$ for studying large scale Physician Referral Networks.
  • Network Specifications. We only build and analyze unweighted, undirected networks, which may overlook the complexities of real-world healthcare interactions. We hope to extend our analyses to more complex network constructions in the future.
  • Feature Completeness. Our prototype may lack some critical features necessary for a comprehensive understanding of healthcare management. This prototype serves as a foundation for future enhancements, and can easily be updated as necessary.
@@ 6.5 Future Work
  • Network Evolution. Studying how healthcare networks evolve over time to identify trends and predict future changes.
  • Thorough Data Analysis. Conducting a more detailed analysis of existing data to derive actionable insights that can influence policy and management decisions.
  • Regional Characteristics. Understanding the distinctive features of regional healthcare networks and how they relate to specific policies and practices.
  • Edge Prediction. Developing methods for predicting new referrals that could alleviate bottlenecks in patient flow, utilizing negative curvature as a predictive metric.
@@ Acknowledgements
We would like to acknowledge the authors and developers of the Datasette package, whose work has been instrumental in the development of our open-source prototype. Their tool has enabled us to efficiently manage, explore, and visualize our dataset, significantly enhancing our ability to analyze and share our findings. We also thank the members of the CARE-CONNECT team, as well as Corinna Coupette for her valuable insights during the early stages of the project and her assistance in creating visualizations, and Katharina Limbeck for her helpful review of the final manuscript. Research reported in this publication was supported by the National Heart, Lung, And Blood Institute of the National Institutes of Health under Award Number R01HL167816. The content is solely the responsibility of the authors and does not necessarily represent the official views of the National Institutes of Health. B.R. is partially supported by the Bavarian state government with funds from the Hightech Agenda Bavaria.